OpenStudy (anonymous):

can someone help me solve √2x -7+8=11

OpenStudy (venomblast):

just pretend the square root are x

OpenStudy (venomblast):

so pretend the original equation is this 5x=x-12

OpenStudy (venomblast):

subtract the x

hartnn (hartnn):

ricardo, just to clarify, is the right side \(\sqrt{x-12}\) or \(\sqrt x-12\) ?

OpenStudy (anonymous):

yes hartn

OpenStudy (venomblast):

set them back to roots \[4\sqrt{x}=-12\]

hartnn (hartnn):

which one ? 1st one or second ?

OpenStudy (venomblast):

divide by 4

OpenStudy (anonymous):

so in other words Venomblast I get rid of x correct

OpenStudy (anonymous):

first

OpenStudy (venomblast):

then square both side to get x alone

OpenStudy (venomblast):